### v1

# Analysis of the Rankine attraction term in an equation-of-state based on the London dispersion force

Abstract
The attraction term in an equation of state for gases, $ac^{2}$, proposed by Rankine in 1854, is generally related to the London dispersion force via a calculation of the second virial coefficient, $B_{2}$, by an equation $B_{2}=2\pi N_{0}\int_{0}^{\infty}\left(1-\exp\left(\omega/kT\right)\right)r^{2}\text{d}r$, where $\omega$ is the potential of the attraction between two molecules in the gas. Here we present an alternative approach that does not use this equation, and does not a-priori assume that the function is quadratic in concentration, $c$. Still, the quadratic dependence on concentration is also found. We analyze the gas argon at temperatures between 300 and 800 K. From the numerical calculations, we derive that the attraction parameter depends on temperature according to a -1/6 power scaling, and thus the attraction component to the second virial coefficient, $B_{2}$, scales with a -7/6 power to temperature. For the same conditions, the virial equation presented above results in a square root scaling of $B_{2}$ with temperature, which is less accurate.
In 1854, Rankine proposed to extend the ideal gas law with an attraction term, $ac^{2}$, where c is concentration, and a an attraction factor. For atoms or molecules described as hard spheres, a correction due to volume exclusion was proposed by Jäger in 1896 as $P\mathrm{{}_{hs}}=cRT\cdot\left(4\eta+10\eta^{2}+\dots\right)$, where $\eta$ is the volume fraction, $\eta=v\mathrm{{}_{m}}c$, with $v\mathrm{{}_{m}}$ the molecular volume (also used by Guggenheim in 1965). The complete equation of state then becomes
$P=P\mathrm{{}_{id}}+P\mathrm{{}_{hs}}+P\mathrm{{}_{att}}=cRT\cdot\left(1+4\eta+10\eta^{2}+\dots\right)-ac^{2}\,.$ (1)
In this paper we focus on the attraction term. Rankine himself included that the factor a is inversely proportional to temperature, i.e., it depends on temperature according to $1/T$. Later, Redlich and Kwong (1949) reduced the dependence of a on temperature to a proportionality with $1/\sqrt{T}$, i.e., $a\propto 1/\sqrt{T}$.
The question is whether the quadratic dependence of the attraction term can can be derived from (i.e., is predicted by) the London dispersion force, and what is then the dependence of a on temperature. The London dispersion force describes that the energy of intermolecular attraction depends on one over the sixth power of distance between the molecules. In literature, the expression for the London dispersion force is then implemented in a function $B_{2}=2\pi N_{0}\int\left(1-\exp\left(\omega/kT\right)\right)r^{2}\text{d}r$ ($\omega<0$) to derive the second virial coefficient, $B_{2}$, which apart from a factor RT equals the attraction parameter a. In this approach, it is assumed that the attraction will only be via a quadratic dependence on c. Assuming that $\omega$ is independent of temperature, then for a high enough temperature, $B_{2}$ has an inverse proportionality on T, and thus the virial equation predicts that the attraction parameter, a, is independent of temperature. But for lower temperatures, the predicted dependence of $B_{2}$ and a on temperature can be different.
We evaluate the energy of a gas due to the London dispersion force (attractive force) as follows. Starting from a central molecule, we evaluate the distribution around it of the first, second, third, etc., closest neighbors, and the energy of their interaction, then multiply this result with concentration, to arrive at an energy density (also divide by 2), and take an appropriate derivative to obtain the contribution to pressure. This analysis is very different from the approach discussed above to derive the second virial coefficient via an integration over space of $\exp\left(\omega/kT\right)$.
The energy of attraction due to London dispersion is
$\omega=-\frac{\lambda}{r^{6}}$ (2)
with $\lambda$ the London dispersion factor, also written as $C\mathrm{{}_{disp}}$, with unit J.m6. It relates to the polarizability of the gas molecules, $\alpha$, and to an energy $h\nu$, according to $\lambda=\tfrac{3}{4}\alpha^{2}h\nu$, where h is the Planck constant and $\nu$ the orbiting frequency of an electron. Dependent on the molecule, this energy is in the range $h\nu=\left\{1.8-3.4\right\}\cdot 10^{-18}$ J. The polarizability of each molecule (atom) is tabulated, such as $\alpha=0.21\penalty 10000\ \AA ^{3}$ for He, and $\alpha=1.64\penalty 10000\ \AA ^{3}$ for Ar. The London dispersion factor, $\lambda$, is independent of temperature. [^1] And for that reason, the Hamaker constant, $A=\pi^{2}q^{2}\lambda$, is temperature-independent as well; here, q is the number density of atoms in a material. In the derivation below, $\lambda$ is equated to the product of kT and a specific length, $r\mathrm{{}_{x}}$, to the power 6, i.e., $\lambda=kT\cdot r_{\text{x}}^{6}$. Because $\lambda$ does not depend on temperature, $r_{\text{x}}^{3}$ is inversely proportional to the square root of temperature, $r_{\text{x}}^{3}\propto 1/\sqrt{T}$.
The energy of attraction of a molecule with its nearest neighbor, due to London dispersion, is
$\Omega=\frac{\int\exp\left(-\omega/kT\right)\cdot\omega\cdot N_{1}\,\text{d}r}{\int\exp\left(-\omega/kT\right)\hphantom{\cdot\omega}\cdot N_{1}\,\text{d}r}$ (3)
where the integration limits go from two times the radius of the atoms, $r=2a$ (assuming the atoms are spherical), to infinity. We implement here a function $N_{1}$, which is the nearest neighbor distribution function, given by
$N_{1}=4\pi r^{2}\cdot n\cdot\exp\left(-\tfrac{4}{3}\pi r^{3}n\right)$ (4)
where n is a concentration in numbers per volume, $n=c\,N_{0}$, with $N_{0}$ the Loschmidt number ($N_{0}=6.022\cdot 10^{23}$ mol-1). [^2] Eq. (4) is based on a derivation that analyzes the distribution of distances from r=0r\!=\!0 onward. That is different in Eq. (3) because we start at r=2ar\!=\!2a. But we can still use it because the $N_{1}$-function is used in the numerator and denominator in the same way, with the same limits of integration.
We can implement Eqs. (2) and (4) in Eq. (3), and arrive at
[…]
Next, we implement a dimensionless, inverse, coordinate, $y=r_{\text{x}}/r$, and then obtain
$\frac{\Omega}{kT}=-\frac{\int\exp\left(y^{6}\right)\cdot y^{2}\hphantom{{}^{-}}\cdot\exp\left(-\delta\cdot y^{-3}\right)\,\text{d}y}{\int\exp\left(y^{6}\right)\cdot y^{-4}\cdot\exp\left(-\delta\cdot y^{-3}\right)\,\text{d}y}$ (6)
where $\delta=\tfrac{4}{3}\pi r_{\text{x}}^{3}n$. Interestingly, the only point where concentration enters the calculation is via $\delta$. Temperature enters the calculation via $r_{\text{x}}$, so also via $\delta$, and in addition via the upper integration limit (which also depends on the atom size).
Eqs. (5) and (6) only consider interaction with the first neighbor, but we can simply add contributions of the second, third, etc., neighbor. We must only replace in each case the function for ${N}_{1}$ by the general expression for the distribution of the ith neighbor, given by
$N_{i}=\frac{3}{\left(i-1\right)!}\left(\tfrac{4}{3}\pi n\right)^{i}r^{3i-1}\exp\left(-\tfrac{4}{3}\pi r^{3}n\right)\,.$ (7)
The generalization of Eq. (6) then becomes
$\frac{\Omega}{kT}=-{\sum}_{i}\left\{\frac{\int\exp\left(y^{6}\right)\cdot y^{\hphantom{-}5-3i}\cdot\exp\left(-\delta\cdot y^{-3}\right)\,\text{d}y}{\int\exp\left(y^{6}\right)\cdot y^{-1-3i}\cdot\exp\left(-\delta\cdot y^{-3}\right)\,\text{d}y}\right\}\,.$ (8)
In a first calculation, we use $\alpha\!=\!1.63\penalty 10000\ \AA ^{3}$ as for Ar, and $h\nu=2.53\cdot 10^{-18}$ J, and thus the London dispersion constant is $\lambda=5.0\cdot 10^{-78}$ J.m6. We calculate $r_{\text{x}}^{3}$ from $r_{\text{x}}^{3}=\sqrt{\lambda/kT}$, and thus at T=300T\!=\!300 K, we have $r_{\text{x}}\!=\!3.27\penalty 10000\ \AA$ . We take a radius $a\!=\!1.72\penalty 10000\ \AA$ , and numerically integrate Eq. (8) from y=0y\!=\!0 (infinitely far) to $y=r_{\text{x}}/\left(2a\right)$, which at 300 K is y=0.950y\!=\!0.950.
After generating calculation output for $\Omega$ as function of $\delta$, thus as function of concentration $n$, we find the pressure according to
$P\mathrm{{}_{att}}=\tfrac{1}{2}\,n^{2}\,\frac{d\Omega}{dn}\,.$ (9)
We evaluate this attractive pressure due to London dispersion, $P\mathrm{{}_{att}}$, in a range of concentrations, c, from 0 to 500 mM, and find a perfect quadratic dependence, with a prefactor a=0.122a=-0.122 J.m3/mol2. Dividing $a$ by RT results in an attraction contribution to the second virial coefficient of $B\mathrm{{}_{2,att}}=-49$ cm3/mol. That is much more negative than an aggregate value for Ar from literature that is $B_{2}=-16$ cm3/mol, but in that number also the hard sphere effect is included. That contribution we estimate as $B\mathrm{{}_{2,hs}}=34$ cm3/mol, and then $B\mathrm{{}_{2,att,lit}}=-50$ cm3/mol, which is close.
We redo the entire calculation at T=400T\!=\!400 K and 500 K. From the reported aggregate values of $B_{2}$, we again subtract $B\mathrm{{}_{2,hs}}$, and thus assume the literature values for attraction are $B\mathrm{{}_{2,att,lit}}=-35$ and 27-27 cm3/mol, respectively. We again find that the contribution of attraction to the pressure scales with concentration squared, and for $B\mathrm{{}_{2,att}}$ we calculate for 400 and 500 K values of $B\mathrm{{}_{2,att}}=-35$ and 27-27 cm3/mol, respectively, in perfect agreement with the experimental data.
The calculation output shows that the attraction factor a scales with temperature to the power $\tfrac{1}{6}$, i.e., $a\propto 1/\sqrt[6]{T}$, and thus our theoretically derived $B\mathrm{{}_{2,att}}$ scales with temperature to the power $\tfrac{7}{6}$. Also calculations at 600 K and 700 K come out close to the data, with the predicted $B\mathrm{{}_{2,att}}$ given by 22-22 and 18-18 cm3/mol, resp., while the literature data for $B\mathrm{{}_{2,att,lit}}$ are 22-22 and 19-19 cm3/mol, resp. At T=200T\!=\!200 K, we calculate $B\mathrm{{}_{2,att}}=-85$ cm3/mol, while $B\mathrm{{}_{2,att,lit}}=-82$ cm3/mol, again close. (Note that at T=200T\!=\!200 K, the attraction parameter comes out 8% larger than the 1/6th scaling.) At T=160T\!=\!160 K, $B\mathrm{{}_{2,att}}=-122$ cm3/mol, while $B\mathrm{{}_{2,att,lit}}=-110$ cm3/mol, thus a deviation develops now that we come closer to the critical point. (At T=160T\!=\!160 K, the attraction parameter is 20% larger than the 1/6th scaling.) Also of interest is that at these temperatures of 200 K and lower, we need to incorporate more and more ‘neighbors’ to make the pressure curve stay quadratic with concentration. Analyzing the first 4 neighbors as we used at higher temperatures, no longer suffices. Our calculated values for $B_{\text{2,att}}$ are plotted in Fig. 1 as continuous line, together with literature data.
Thus, we propose for Ar in the temperature range 300800300-800 K the equation of state
$P=cRT\cdot\left(1+4\eta+10\eta^{2}+20\eta^{3}\right)-a^{*}\cdot T^{-1/6}\cdot c^{2}$ (10)
[…]
In the Redlich-Kwong theory, the attraction term is modified by dividing by $1+bc$, where $b=4v\mathrm{{}_{m}}$. We implement this term in Eq. (10), resulting in
$P=cRT\cdot\left(1+4\eta+10\eta^{2}+20\eta^{3}\right)-a^{*}\cdot T^{-1/6}\cdot c^{2}/\left(1+\alpha\eta\right)$ (11)
which gives an improved description of the critical point, especially when we allow the factor $\alpha$ not to be 4, but to be reduced to $\alpha=1.2$. Then the equation predicts $P\mathrm{{}_{c}}=5.5$ MPa, $T\mathrm{{}_{c}}=150$ K, and $c\mathrm{{}_{c}}=13.2$ M, which is very close to the data. Thus, small modifications to the attraction term, that at low concentrations do not have an effect, can markedly modify the prediction of the critical point. In this case, the attraction term is only modified by 10% at the critical point.
Using the virial approach, $B_{2}=2\pi N_{0}\int\left(1-\exp\left(\omega/kT\right)\right)r^{2}\text{d}r$ ($\omega<0$), which we solve using the same atomic radius, a, and same $\lambda$ as above, we obtain $B_{2}=-72$ cm3/mol at 300 K, 61-61 cm3/mol at 400 K, and 54-54 cm3/mol at 500 K. These calculated numbers for $B_{2}$ underestimate the influence of temperature on $B_{2}$, because they scale approximately with $1/\sqrt{T}$. And thus the interaction parameter a is predicted to scale with $\sqrt{T}$, i.e., increases with temperature, which is not so likely.
In conclusion, we presented an analysis of the London dispersion force based on a method that considers interactions of the first, second, third, etc., neighbors around a molecule, and for argon in the temperature range 300800300-800 K, we find a scaling of the attraction parameter, a, in the Rankine attraction term, $ac^{2}$, that is proportional to temperature to the power 1/6-1/6. This result is in agreement with data for $B_{2}$ of argon in that temperature range.
Figure 1: The contribution of attraction to the second virial coefficient of argon, as function of temperature. Continuous line is based on evaluation of the London dispersion force by the method explained in this paper, and circles are literature data. Dashed line is a 7/6th scaling with temperature. These data are obtained from aggregate $B_{2}$-values from which we subtract $B_{\text{2,hs}}=34$ cm3/mol, which is the contribution of hard sphere, volumetric, interactions.